Given the dataset in the table below:

10

Beer, Nuts, Diapers

20

Beer, Coffee, Diapers

30

Beer, Diapers, Eggs

40

Nuts, Eggs, Milk

50

Beer, Coffee, Milk

60

Diapers, Eggs, Milk

70

Coffee, Diapers, Eggs

80

Beer, Nuts, Coffee, Diapers, Eggs, Milk

What is the support and confidence of the rule: Nuts -> Beer?

Group of answer choices

1. [support=37.5%, confidence=60%]

2. [support=25%, confidence=40%]

3. [support=37.5%, confidence=75%]

4. [support=25%, confidence=66.66%]
Answer #1

Rule is Nuts -> Beer

A = Nuts

B = Beer

Support = Transactions containing both A and B / Total number of transactions

Support = 2 / 8 = 0.25

Support % = 0.25 * 100 = 25%

Confidence = Transactions containing both A and B / Transactions containing A

Confidence = 2 / 3 = 0.6666

Confidence % = 0.6666 * 100 = 66.66%

4th option [support=25%, confidence=66.66%] is correct.